milana:

Select the number sentence that illustrates the associative property of multiplication. * A. (4 × 6) × 7 = 4 × (6 × 7) B. (4 × 6) × 7 = 7 × (4 × 6) C. 4 × 6 = 6 × 4 D. 4 × 6 = 7 × (6 × 4) answer the correct

Extrinix:

Associative Property: \(a+(b+c)=(a+b)+c\) Property of Multiplication: \(a\times b=b\times a\) So combined would look like this: \(a\times (b\times c)=(a\times b)\times c\)

milana:

not u go i just need the answer!!!!